### v1

# DR-RAG: Applying Dynamic Document Relevance to Retrieval-Augmented Generation for Question-Answering

Abstract
Retrieval-Augmented Generation (RAG) has significantly demonstrated the performance of Large Language Models (LLMs) in the knowledge-intensive tasks, such as Question-Answering (QA). RAG expands the query context by incorporating external knowledge bases to enhance the response accuracy. However, it would be inefficient to access LLMs multiple times for each query and unreliable to retrieve all the relevant documents by a single query. We find that even though there is low relevance between some critical documents and query, it is possible to retrieve the remaining documents by combining parts of the documents with the query. To mine the relevance, a two-stage retrieval framework called Dynamic-Relevant Retrieval-Augmented Generation (DR-RAG) is proposed to improve document retrieval recall and the accuracy of answers while maintaining efficiency. Also, a small classifier is applied to two different selection strategies to determine the contribution of the retrieved documents to answering the query and retrieve the relatively relevant documents. Meanwhile, DR-RAG call the LLMs only once, which significantly improves the efficiency of the experiment. The experimental results on multi-hop QA datasets show that DR-RAG can significantly improve the accuracy of the answers and achieve new progress in QA systems.
DR-RAG: Applying Dynamic Document Relevance to Retrieval-Augmented Generation for Question-Answering
Zijian Hei∗,1,2   Weiling Liu∗,1,3   Wenjie Ou∗,1,4   Juyi Qiao1   Junming Jiao1   Zhiqing Zhu1   Guowen Song†,1 1Li Auto Inc.   2Sun Yat-sen University   3Northeastern University, China   4Sichuan University thanks:  Equal contribution to this work during internshiping at Li Auto.    † Corresponding author 1 {songguowen, zhuzhiqing}@lixiang.com
1 Introduction
Large language models (LLMs) have recently made significant improvement in the field of Natural Language Processing (NLP), especially in text generation tasks Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023b); Anil et al. (2023); Ouyang et al. (2022); Touvron et al. (2023a). Although LLMs perform well in various application scenarios, there are still challenges to the accuracy and timeliness of the generated text, such as in real-time domains. LLMs with intrinsic parameter memories may generate inaccurate or even incorrect text when faced with up-to-date query Min et al. (2023); Mallen et al. (2022); Muhlgay et al. (2023). The problem is known as hallucination, where the text generated by LLMs fail to correspond to real-world knowledge Ji et al. (2023); Zhang et al. (2023); Kwiatkowski et al. (2019). Therefore, Retrieval-Augmented Generation (RAG) frameworks are proposed to improve the accuracy of generated text by combining relevant information from external knowledge base with query Arora et al. (2023); Lewis et al. (2020); Borgeaud et al. (2022). RAG has effectively demonstrated its superiority in knowledge-intensive tasks such as open-domain Question-Answering (QA) and has achieved new progress in the LLMs’ performance.
Figure 1: QA performance (F1) and time for different RAG frameworks. We use the GPT-3.5-Turbo-Instruct as the base LLM on the multi-hop QA datasets (MuSiQue, HotpotQA and 2Wiki).
However, irrelevant information reduces the quality of the generated text and further interferes with the ability of LLMs to answer the query in the application Shi et al. (2023). Moreover, the undifferentiated combining strategy in RAG can lead to mixing in some irrelevant information Rony et al. (2022). Inconsistent or contradictory information during combining the document may lead to the introduction of incorrect information and have an impact on the accuracy of the generated answers. In the retrieval, we need to select documents that are highly relevant and decisive for the generation of answers (static-relevant documents) and documents that are low relevant but also crucial to the generation of answers (dynamic-relevant documents). As shown in Fig. 2, an example query is ‘Who is the spouse of the child of Peter Andreas Heiberg?’, which requires the two most relevant documents to obtain the correct answers. Static-relevant documents is easy to be retrieved due to the high relevance with the query on ‘Peter Andreas Heiberg’ and ‘child/son’ (Fig. 2 red). However, dynamic-relevant documents is difficult to be retrieved because it is only related to the query as a ‘spouse/wife’ (Fig. 2 blue). Moreover, the knowledge base contains too much information about ‘spouse’, which may cause dynamic-relevant documents to be ranked lower in the retrieval process. There is a high relevance on ‘Johan Ludvig Heiberg’ and ‘wife’ between static- and dynamic-relevant documents. If ‘spouse/wife’ with the query is also taken into account, we can easily retrieve dynamic-relevant documents to get the answer.
Figure 2: An example shows that retriever easily introduces static-relevant documents due to high relevance (red), but struggles to retrieve dynamic-relevant documents which are of low relevance (blue) but critical for the answer. Stars are levels of retrieval difficulty.
Motivated by the above observations, a novel two-stage retrieval framework called Dynamic-Relevant Retrieval-Augmented Generation (DR-RAG) is proposed to mine the relevance between the query and documents. In the first-retrieval stage, similarity matching (SM) method is used to obtain a certain percentage of documents based on the query. Subsequently, the documents with the query are concatenated to dig further into more in-depth relevance to dynamic-relevant documents. Moreover, we design a classifier that determines whether the retrieved documents contribute to the current query by a predefined threshold. To optimise the documents, we design two approaches, i.e., forward selection and reverse selection. We aim to ensure that the retrieved documents are highly relevant, thus avoiding redundant documents. Through two-stage retrieval and classifier selection strategies, DR-RAG has the ability to retrieve sufficient relevant documents and address complex and multilevel problems. DR-RAG can make full use of the static and dynamic relevance of documents and enhance the model’s performance under diverse queries. To validate the effectiveness of DR-RAG, we conduct extensive experiments by different retrieval strategies on multi-hop QA datasets. The results show that our method can significantly improve the recall and accuracy of the answers.
In short, we summarize the key contributions of this work as follows:
  • We design an effective RAG framework named DR-RAG, which is effective in multi-hop QA. Two-stage retrieval strategy is proposed to significantly improve the recall and accuracy of the retrieval results.
  • We design a classifier that determines whether the retrieved documents contribute to the current query by setting a predefined threshold. The mechanism can effectively reduces redundant documents and ensures that the retrieved documents are concise and efficient.
  • We conduct experiments on three multi-hop QA datasets to validate our DR-RAG. The experimental results show that DR-RAG has the ability to improve recall by 86.75% and improve by 6.17%, 7.34%, 9.36% in the three metrics (Acc, EM, F1). DR-RAG has significant advantages in complex and multi-hop QA and support the performance of the RAG frameworks in QA systems.
2 Related Works
2.1 RAG for Multi-hop QA
RAG is a popular framework for LLMs and has received much attention to many tasks, such as QA systems. RAG Lewis et al. (2020) combined a sequence-to-sequence model with external knowledge bases to significantly improve the quality of quizzing and summarization tasks. The decomposition of a complex query Khattab et al. (2022); Press et al. (2022); Pereira et al. (2023); Khot et al. (2022); Zhou et al. (2022) into a series of simpler sub-queries might inevitably require multiple calls to LLMs, resulting in high computational cost. Adaptive-RAG Jeong et al. (2024) evaluated the complexity of the problem by a classifier and selects the most appropriate retrieval strategy based on the classification results. RQ-RAG Chan et al. (2024) aimed to improve the performance of models by optimising search query, including rewriting, decomposition and disambiguation. However, it would be inefficient to access LLMs multiple times for each query and unreliable to retrieve all dynamic-relevant documents by a single query.
Table 1: The key mathematical notations.
Notation Description
$\boldsymbol{q}$ the user’s input query
$\boldsymbol{q}^{*}$ the query combined with retrieved document after the first-retrieval stage
$D$ the knowledge base for storing documents
$C$ the trained classifier
$k$ the total number of documents retrieved from $D$
$k_{1}$ the total number of documents retrieved from $D$ in the first-retrieval stage
$k_{2}$ the total number of documents retrieved from $D$ in the second-retrieval stage
$n$ the number of documents critical for correctly answering $\boldsymbol{q}$
$\boldsymbol{d}$ the documents retrieved from $D$
$\boldsymbol{d}^{*}$ the relevant documents for correctly answering $\boldsymbol{q}$
$\boldsymbol{d}^{\Delta}$ the irrelevant documents for correctly answering $\boldsymbol{q}$
$\boldsymbol{d_{stat}^{*}}$ the documents with static relevance to the query
$\boldsymbol{d_{dyn}^{*}}$ the documents with dynamic relevance to the query
2.2 Retriever in RAG
The retriever in a RAG system is the key to verify how the retriever can obtain revelant instant contexts from external knowledge bases and alleviate the hallucination of LLMs. Fan et al. (2021) combined $K$ Nearest Neighbor (KNN) retrieval with a traditional transformer model to dynamically access historical data and provide enough information by a composite memory. Cheng et al. (2024) proposed Selfmem to make the generated text more relevant to the retrieved information through a self-memory mechanism. Recent research has highlighted the potential applications of LLMs, which can be considered as supervised signals for training retrieval components, even as retrieval components. These findings provide us with new avenues for exploring the ability of retrievers to improve the efficiency of information retrieval based on the document relevance. In our work, we retrieve multiple relevant documents based on the query by a two-stage strategy and design a classifier to determine whether the documents can answer the query, and the remaining relevant documents are fed into LLMs with the query to obtain the answer.
3 Method
Figure 3: An overview of DR-RAG. In step 1, we retrieve static-relevant documents (SR-Documents) due to high relevance with the query. Then we concatenate SR-Documents with the query to retrieve multiple dynamic-relevant documents (DR-Documents) in step 2. Finally, we select each of DR-Documents in turn to splice with the query and SR-Documents and feed them into the classifier to select the most relevant DR-Document.
In this section, we will describe the DR-RAG framework and its design approach in detail. Specifically, in section 3.1 we will define relevant symbols comprehensively, and in section 3.2 we will describe the whole framework.
3.1 Preliminaries
To enrich the knowledge of LLMs, we need to retrieve multiple documents to provide comprehensive answers to complex query. For better clarity, we summarize the key notations in Table 1 and the whole framework can be referred to in Fig. 3.
Our goal is to retrieve the most relevant documents $\boldsymbol{d}^{*}$ from the retrieved documents $\boldsymbol{d}$ to answer the query and prevent missing key information from the additional knowledge provided to LLMs. However, it is difficult to retrieve all the static- and dynamic-relevant documents through SM method during the retrieval process (Fig. 3). For clearness, we name these two types of relevant documents as $\boldsymbol{d_{stat}^{*}}$ and $\boldsymbol{d_{dyn}^{*}}$, respectively.
A common approach is to increase the value of $k$ to expand the possibility of retrieving $\boldsymbol{d_{dyn}^{*}}$. For instance, in MuSiQue, increasing $k$ from 3 to 6 only raises the recall rate from 58% to 76%, leaving many relevant documents unretrieved. Furthermore, irrelevant documents will provide LLMs with redundant information. Motivated by the problem, the main research objective of our work is to improve the document recall rate of $\boldsymbol{d_{dyn}^{*}}$ based on dynamic relevance with the same top-$k$.
3.2 DR-RAG
In this section, we will give a comprehensive description about the DR-RAG framework, a new two-stage retrieval method compared to traditional reranking methods NetEase Youdao (2023); Chen et al. (2024). From Fig. 3, we retrieve $k_{1}$ documents through SM method (first-retrieval stage) and employ a classifier $C$ to model the dynamic relevance between documents (selection process) to enhance the recall rate of the remaining $k_{2}$ documents. The classifier $C$ lies in assessing the dynamic relevance between documents to determine whether the information from the documents is crucial to answer the present query. DR-RAG aims to retrieve $\boldsymbol{d_{stat}^{*}}$ from the $k_{1}$ documents in the first-retrieval stage and then retrieve $\boldsymbol{d_{dyn}^{*}}$ from the $k_{2}$ documents (second-retrieval stage). By combining these key documents, LLMs have the ability to generate the answer of queries more accurately.
3.2.1 Query Documents Concatenation
As mentioned before, due to the low relevancy between dynamic-relevant documents and the query, the documents are difficult to be retrieved. Moreover, the only relevant information ‘spouse/wife’ between them is also obscured by the mixed information in the knowledge base because too many documents in $D$ will contain ‘spouse’. Therefore, Query Documents Concatenation (QDC) method aims to employ the sentence to match for more useful and relevant information. After the first-retrieval stage, we will obtain $k_{1}$ static-relevant documents and concatenate $\boldsymbol{q}$ with each document to form multiple $<\boldsymbol{q},\boldsymbol{d_{i}}>,i\in k_{1}$ pairs. Moreover, dynamic-relevant documents from $D$ can be retrieved by corresponding $<\boldsymbol{q},\boldsymbol{d_{i}}>,i\in k_{1}$ pair in the second-retrieval stage. As the case in Fig 3, when $\boldsymbol{q}$ and $\boldsymbol{d_{stat}^{*}}$ are concatenated, the query contains both the name ‘Johan Ludvig Heiberg’ and the relationship ‘spouse/wife’, which is essentially similar to $\boldsymbol{d_{dyn}^{*}}$. Therefore, $\boldsymbol{d_{dyn}^{*}}$ is more clearly related to the query and thus easily retrieved. The whole process is:
$\begin{split}\boldsymbol{d_{m}}&=\texttt{Retriever}(\boldsymbol{q}),m\in[1,k_{1}]\\ \boldsymbol{q}^{*}_{m}&=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d_{m}})\\ \boldsymbol{d_{n}}&=\texttt{Retriever}(\boldsymbol{q}^{*}_{m}),n\in[k_{1}+1,k_{1}+k_{2}]\\ answer&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{d_{1}},\boldsymbol{d_{2}},...\boldsymbol{d_{k_{1}+k_{2}}},\boldsymbol{q}))\\ \end{split}$ (1)
where $k_{1}+k_{2}$ is equal to $k$. Retriever is a common SM method. $\boldsymbol{d_{i}},i\in[1,k]$ is the $i$-th relevant document retrieved from $D$. LLM is a large language model to obtain the answer. $answer$ represents the output to answer the query.
3.2.2 Classifier for Selection
While QDC method significantly improves document recall and answer accuracy, there are two key issues to consider: 1) There may be redundant information in the $k$ retrieved documents, which may affect the response of LLMs; 2) How to determine whether a document retrieved in the second-retrieval stage is valid for an answer to further optimise document recall. Motivated by the issues, two pipelines are designed to dig into in-depth document relevance and solve the issues: 1) Classifier Inverse Selection (CIS): in this pipeline, after the second-retrieval stage we exclude some irrelevant documents from the $k$ retrieved documents; 2) Classifier Forward Selection (CFS) : we set a judgment condition to each retrieved document in the second-retrieval stage to filter out irrelevant documents which are useless or even play a negative role in the answer. In addition, we will train a classifier $C$ by a small model with millisecond-level runtime to prevent excessive delays in our pipelines. DR-RAG involves a small binary-classification model where the input consists of $\boldsymbol{q}$ and two documents. The training objective is to determine the potential contribution of the documents to answering $\boldsymbol{q}$. The specific settings are as follows:
$\begin{split}\texttt{Classifier}(\boldsymbol{q},\boldsymbol{d}^{*},\boldsymbol{d}^{*})&=positive\\ \texttt{Classifier}(\boldsymbol{q},\boldsymbol{d}^{*},\boldsymbol{d}^{\Delta})&=negative\\ \texttt{Classifier}(\boldsymbol{q},\boldsymbol{d}^{\Delta},\boldsymbol{d}^{\Delta})&=negative\\ \end{split}$ (2)
where Classifier represents the binary classification model. $positive$ and $negative$ results indicate whether the two documents are critical for the query.
Classifier Inverse Selection    In this approach, we selectively exclude some irrelevant documents from the retrieved $k$ documents to minimize document redundancy. Specifically, after obtaining $k$ documents in stages, we pair them pairwise as $<\boldsymbol{q},\boldsymbol{d_{m}},\boldsymbol{d_{n}}>$ and get $C_{k}^{2}$ pairs. The pairs, with the current query $\boldsymbol{q}$, are collectively fed into the classifier $C$. Similarly, when the classification result of a document and the remaining $k$-1 documents is negative, then we consider the document as redundant and should be removed. The whole process is:
Table 2: Results on different datasets with Llama3 (8B) as LLM. We emphasize our results in bold.
MuSiQue HotpotQA 2Wiki
Methods EM F1 Acc Step Time EM F1 Acc Step Time EM F1 Acc Step Time
Single-step Approach 13.80 22.80 15.20 1.00 1.00 34.40 46.15 36.40 1.00 1.00 41.60 47.90 42.80 1.00 1.00
Adaptive Retrieval 6.40 15.80 8.00 0.50 0.55 23.60 32.22 25.00 0.50 0.55 33.20 39.44 34.20 0.50 0.55
Self-RAG 1.60 8.10 12.00 0.73 0.51 6.80 17.53 29.60 0.73 0.45 4.60 19.59 38.80 0.93 0.49
Adaptive-RAG 23.60 31.80 26.00 3.22 6.61 42.00 53.82 44.40 3.55 5.99 40.60 49.75 44.73 2.63 4.68
Multi-step Approach 23.00 31.90 25.80 3.60 7.58 44.60 56.54 47.00 5.53 9.38 49.60 58.85 55.40 4.17 7.37
DR-RAG(Ours) 26.97 38.90 34.03 1.00 1.54 48.58 62.87 55.68 1.00 1.40 49.60 55.62 55.18 1.00 1.43
Classifier Forward Selection    Unlike the CIS method proposed above, CFS method aims to remove the irrelevant dynamic-relevant documents in the second-retrieval stage. To achieve this goal, we search for a document $d_{step2}$ from $D$ according to the $<\boldsymbol{q},d_{step1}>$ pair, and feed both the query and documents into $C$. When the classification result is negative, we will exclude the dynamic-relevant document in the current retrieved documents, and search for the next dynamic-relevant document which can be classified as positive with $d_{step1}$.
4 Experiment Settings
The experimental details will be described in this section. Due to space constraints, the descriptions of implementation details, retrieval strategy and baseline can seen in Appendix A.2, A.3 and A.4.
4.1 Dataset
We verify the effectiveness of our proposed framework on three multi-hop QA datasets, including HotpotQA, 2Wiki and MuSiQue Yang et al. (2018); Ho et al. (2020); Trivedi et al. (2022b). The datasets require the system to comprehensively collect and contextualize information from multiple documents to answer more complex queries.
Table 3: Results on different LLMs compared to Adaptive-RAG. We emphasize our best results in bold.
MuSiQue HotpotQA 2Wiki
top-$k$ LLMs Methods EM F1 Acc EM F1 Acc EM F1 Acc
Adaptive-RAG 23.60 31.80 26.00 42.00 53.82 44.00 40.60 49.75 46.40
3 gpt-3.5 Query Document Concatenation 21.40 32.20 29.90 37.80 51.56 52.60 36.20 48.99 51.40
Classifier Inverse Selection 23.70 33.70 29.40 41.20 53.91 53.40 38.00 51.48 54.20
Classifier Forward Selection 26.00 36.20 35.00 43.80 58.83 55.00 48.40 60.13 64.20
Llama3-8B Query Document Concatenation 21.30 32.00 27.10 43.85 56.38 49.88 40.84 48.61 43.76
Classifier Inverse Selection 20.90 31.90 27.00 44.88 57.05 50.94 42.71 50.82 46.45
Classifier Forward Selection 26.50 37.40 32.60 48.71 62.42 55.26 50.48 59.29 56.07
4 gpt-3.5 Query Document Concatenation 25.37 25.05 35.70 42.15 55.79 53.31 50.60 59.99 62.20
Classifier Inverse Selection 25.80 36.50 35.60 42.00 56.10 54.36 49.40 60.40 65.00
Classifier Forward Selection 25.80 37.60 38.60 45.00 60.55 57.40 52.40 63.95 69.60
Llama3-8B Query Document Concatenation 25.10 37.10 32.00 46.05 59.98 53.09 45.79 54.62 39.70
Classifier Inverse Selection 25.70 37.50 32.70 48.02 61.43 54.36 50.52 59.54 55.94
Classifier Forward Selection 27.10 39.30 34.30 48.30 62.81 55.22 50.30 59.40 55.85
6 gpt-3.5 Query Document Concatenation 28.80 41.30 38.60 45.41 60.09 61.60 48.20 62.58 67.00
Classifier Inverse Selection 31.20 42.70 40.20 45.80 60.38 60.80 52.60 65.93 71.20
Classifier Forward Selection 28.40 41.10 40.60 48.20 63.83 63.60 49.80 63.74 68.40
Llama3-8B Query Document Concatenation 25.90 38.10 33.70 46.19 60.21 54.14 44.02 53.08 48.73
Classifier Inverse Selection 27.50 39.40 34.60 47.34 61.30 54.55 50.59 59.58 56.29
Classifier Forward Selection 27.30 40.00 35.20 48.73 63.38 56.56 48.02 57.17 53.63
5 Results and Analysis
Table 4: Results of different retrieval strategies. Actual numbers represents the actual document numbers that we feed into LLMs.
MuSiQue HotpotQA 2Wiki
top-$k$ Retrieval Strategies Recall rate Actual numbers Recall rate Actual numbers Recall rate Actual numbers
3 BM25 37.57 3.00 64.67 3.00 57.46 3.00
Similarity Matching 58.31 3.00 80.33 3.00 74.34 3.00
Query Document Concatenation 58.44 3.00 86.12 3.00 74.48 3.00
Classifier Inverse Selection 57.42 2.82 79.80 2.74 74.34 2.60
Classifier Forward Selection 66.45 2.95 88.89 2.89 87.81 2.83
4 BM25 43.52 4.00 70.72 4.00 63.95 4.00
Similarity Matching 66.20 4.00 85.57 4.00 80.05 4.00
Query Document Concatenation 70.30 4.00 90.20 4.00 89.89 4.00
Classifier Inverse Selection 66.45 3.61 88.37 3.43 89.71 3.03
Classifier Forward Selection 73.83 3.82 92.28 3.75 94.13 3.41
6 BM25 51.69 6.00 79.93 6.00 73.95 6.00
Similarity Matching 76.78 6.00 92.59 6.00 88.83 6.00
Query Document Concatenation 79.46 6.00 94.00 6.00 95.04 6.00
Classifier Inverse Selection 77.13 5.23 93.24 4.72 94.79 3.72
Classifier Forward Selection 83.01 5.69 96.27 5.35 98.04 4.63
5.1 Main Results
Experimental results effectively demonstrate the effectiveness of our proposed DR-RAG. Table 2 and 3 presents the performance of DR-RAG in answering multihop query, highlighting the advantages of our approach compared to the current RAG framework Jeong et al. (2024); Asai et al. (2024) across multiple metrics. Table 4 shows the performance of DR-RAG across various retrieval strategies.
As shown in Table 2, DR-RAG achieves higher accuracy across multiple datasets compared to other RAG frameworks, which is in line with our expectations. When retrieving the same $k$ documents, DR-RAG can achieve a higher recall rate and a higher percentage of correct answers. The core idea of our work is to provide enough information for LLMs to answer the query. To fully validate the effectiveness of DR-RAG in multiple dimensions, we follow the metrics (F1, EM, Acc) of Adaptive-RAG and choose the same metrics for the evaluation. From the results, DR-RAG achieves better performance than other baseline RAG frameworks (self-RAG and Adaptive-RAG) on all three metrics. Moreover, DR-RAG is also less than other RAG frameworks in terms of the number of LLMs responses and the time consumed in QA systems.
5.2 Analysis
Effects of Recall Rate    The ability of LLMs to answer domain-specific query correctly almost depends on whether all the necessary information is included in the prompt context. When relevant information is missing, it is difficult for LLMs with the hallucination problem to accurately answer the query. Without relevant information, LLMs may rely on existing fixed knowledge to response the queries and lead to incorrect answers. Therefore, we devise a method to try to improve the retrieval capability and retrieve as enough relevant documents as possible. Table 5 illustrates the answers of the query with and without sufficient information provided to LLMs. As seen in Table 4, in 2Wiki, our retrieval strategy already achieves a recall rate of 98% when top-$k$ is 6. When we feed enough relevant information into LLMs, the accuracy of their answers can be improved accordingly. CFS method achieves higher recall rate by 26.4% and 8.6% than BM25 and SM methods, respectively, which proves the feasibility of DR-RAG.
Table 5: Case study with Llama3-8B, where we present the factual error in red and the accurate infomation in blue.
Datasets Queries Traditional RAG DR-RAG (Ours)
MuSiQue
Who is the spouse of
the Green performer?
Retrieval strategy: SM Reasoning: The spouse of the Green performer is Debbie Allen because she is the sister of the Green perfor- mer’s sister. The answer is: Debbie Allen. Retrieval strategy: CFS Reasoning: The Green performer mentioned in the question is Steve Hillage. Then we i- dentify that Miquette Giraudy is mentioned as his partner. The answer is: Miquette Giraudy.
HotpotQA
what year was the cover
artist of Multiverse: Ex-
ploring Poul Anderson’s
Worlds born ?
Retrieval strategy: SM
Reasoning: The book “Multiverse:
Exploring Poul Anderson’s Worlds”
was first published in hardcover in
May 2014 by Subterranean Press.
And assuming Bob Eggleton was
around 50 years old. We can cal-
culate his birth year is 1964.
The answer is: 1964.
Retrieval strategy: CFS
Reasoning: the cover artist of “Multiverse:
Exploring Poul Anderson’s Worlds” is Bob
Eggleton. We also know that Bob Eggleton
was born on September 13, 1960.
The answer is: 1960
2Wiki
What nationality is the
director of film Downpour
(Film)?
Retrieval strategy: SM
Reasoning: we know that “Down-
pour” is directed by Bahram Bayzai.
the nationality of Bahram Bayzai is
not explicitly mentioned in the pro-
vided contexts. However, based on
my knowledge, Bahram Bayzai is
an Iranian director.
The answer is: Iranian
Retrieval strategy: CFS
Reasoning: It is mentioned that the film
“Downpour” is directed by Bahram Bayzai.
And Bahram Beyzaie is associated with
persian letters, arts. So the director of the
film “Downpour” is of persia nationality.
The answer is persia.
Effects of Redundant Information    We hypothesise that if there is less redundant information in the contextual knowledge, LLMs can fully understand the query to reduce the hallucination. Therefore, CIS method is devised to validate this hypothesis. Invalid information may increase by about 30% as the number of documents fed into LLMs increases, but LLMs fail to judge the information when answering. LLMs may refer to redundant information and provide an answer with incorrect information. The results all validate our hypothesis that we should provide LLMs as little redundant or incorrect information as possible throughout the RAG process. CIS method is effective in removing redundant information, but it may reduce the quality of responses when the reduction in recall is too large. Even though we feed all the relevant documents into LLMs, it is still possible to fail to get the right answer. In Table 4, on dataset 2Wiki, when the number of documents $k$ provided to LLMs at 4 and 6, there is only a slight increase from CIS to CFS in the recall and instead a decrease in the metrics. Therefore, CFS method is propsed to balance redundant and relevant information.
Increase Recall with Lower Documents    In CFS method, it seems impossible to find a match for every $<\boldsymbol{q},\boldsymbol{d}>$ pair in the second-retrieval stage because the documents we need have been retrieved. Therefore, there will be cases where the total number of our retrieved documents is less than $k$. For instance, in the HotpotQA dataset, when $k$ is set to 6, the average number of documents actually provided to LLMs is only 5.35, thereby reducing irrelevant information to some extent. CFS method in Table 4 achieves a higher recall rate while retrieving fewer actual numbers of documents compared to QDC method. CFS method yields higher scores across the three metrics in our experiments and achieves more significant retrieval capabilities with lower redundant inputs than other methods.
Ablation Study
Table 6: Ablation study on MuSiQue by Llama3-8B.
top-k LLMs EM F1 Acc
3 CFS 48.71 62.42 55.26
w/o QDC 46.26 59.37 52.60
CIS 44.88 57.05 50.94
w/o QDC 44.82 57.00 50.88
4 CFS 48.30 62.81 55.22
w/o QDC 47.78 61.63 54.73
CIS 48.02 61.43 54.36
w/o QDC 46.54 59.28 52.80
6 CFS 48.73 63.38 56.56
w/o QDC 48.00 63.10 56.29
CIS 47.34 61.30 54.55
w/o QDC 46.30 60.00 53.95
Time for One Response    Compared to previous RAG frameworks, DR-RAG also achieves better time optimization during the whole process. Other RAG frameworks may call LLMs multiple times before obtaining the answer, resulting in high computational cost. In fact, the inference time of LLMs is also a worthwhile optimization in the applications. It takes a lot of time to call LLMs once, and calling them multiple times presents a catastrophic challenge in terms of time overhead. Therefore, we attempt to design a small model with relatively few parameters to achieve better optimization rather than calling LLMs multiple times. In Table 2, compared to Adaptive-RAG, we have achieved an average 74.2% reduction in time overheads. Therefore, we can conclude that we can achieve better experimental efficiency and the time overhead makes DR-RAG valuable for practical engineering applications.
Case Study    We conduct a case study to qualitatively compare our DR-RAG against the traditional RAG. Table 5 demonstrates the specific inference cases on the multi-hop datasets. For example, in MuSiQue dataset, our DR-RAG identifies the answer to the query by only using the LLM’s parametric knowledge about ‘partner’. Traditional RAG sometimes generate incorrect responses due to the inclusion of irrelevant information about ‘sister’. Meanwhile, faced with a complex query, DR-RAG can first retrieve static-relevant documents based on ‘cover artist’ and ‘Multiverse: Exploring Poul Anderson’s Worlds’ to get the name ‘Bob Eggleton’. Then, in the second-retrieval stage, by combining the name ‘Bob Eggleton’ with ‘born’ in the query, dynamic-relevant documents can be retrieved to obtain the answer ‘1960’.
Optimization under Resource Constraints    The classifier $C$ of document relevance requires certain hardware conditions and resources for data annotation. However, QDC method indicates that dynamic documents relevance can still be utilized without $C$. As seen in Table 4, compared to SM method, across all datasets, when top-$k$ is 4 or 6, there is a significant increase of the retrieval recall by 3.84%. Yet, when top-$k$ is 3, there is also 6% increase on HotpotQA and a slight increase on the other two datasets. This suggests that by making reasonable choices about top-$k$, even in cases where resources are limited, the performance of retrieval can be optimized by leveraging the relevance of relevant documents, thus improving LLMs’ performance in QA tasks.
6 Conclusion
This paper presents DR-RAG, an innovative RAG framework designed to enhance document retrieval accuracy by leveraging the relevance of different documents in various QA scenarios. Throughout this research, we explore diverse retrieval strategies and conduct comprehensive experimental comparisons. Ultimately, we adopt CFS as the final framework, which not only reduces the number of redundant document but also achieves the most superior performance. Additionally, we analyze the utilization of dynamic document relevance under constrained training resources. The experimental results demonstrate that DR-RAG significantly improves answer quality and reduces the time required for QA systems.
Appendix A Appendix
A.1 Limitations
While DR-RAG has demonstrated excellent performance across multiple datasets for multi-hop QA, its implementation requires the prior training of a distinct classifier. It is uncertain whether our classifier will be effective with niche domains. Therefore, DR-RAG can serve as an invaluable inspiration to train a classifier with private data. In the future, we will collect more comprehensive data to train a more applicable classifier for various QA tasks.
A.2 Implementation Details
We follow the standard evaluation approach Jeong et al. (2024) and validate our DR-RAG for QA systems by multiple metrics including F1, EM, and Accuracy (Acc). These metrics provide an objective measure between the prediction results and ground truth. In addition, the efficiency is also another issue we have to tackle. Most existing RAG frameworks Asai et al. (2024); Jeong et al. (2024) require multiple calls to LLMs for inference. Therefore, we consider the number of inferences by LLMs and the time required for responses as our evaluation. To eliminate the effects of different LLMs, we select gpt-3.5-turbo Achiam et al. (2023); Brown et al. (2020) and Llama3-8B Liu et al. (2024) as base LLMs, and accurately acquire the answers to query based on retrieval documents. For the classifier $C$, we fine-tune bigbird-roberta-base Zaheer et al. (2021) by the entire training set to accommodate longer input tokens. Due to the imbalance between positive and negative samples in the datasets, we sample the positive and negative examples to construct the datasets with the ratio of 1:1. In addition, we sample about 2300 pieces of data in each dataset, which exceeds the existing experiment Jeong et al. (2024) in sample numbers.
A.3 Retrieval Strategy
DR-RAG aims to solve the problem of low recall in document retrieval. Therefore, five different retrieval strategies are designed to verify the effectiveness of our proposed DR-RAG.
  • BM25: A method to measure the relevance between $\boldsymbol{q}$ and the documents.
  • SM: The retrieval documents will be embedded and stored in $D$ and the similarity between $\boldsymbol{q}$ and $D$ is calculated to extract the $k$ most relevant documents.
  • QDC: We first retrieve $k_{1}$ documents from $D$, concatenate $\boldsymbol{q}$ with the documents to form multiple pairs and retrieve the $k_{2}$ most relevant documents for $<\boldsymbol{q},\boldsymbol{d}>$ pairs, respectively, until the number of retrieved documents = $k$.
  • CIS: To minimize document redundancy in QDC, all $k$ documents retrieved are pairwise combined, concatenated with $\boldsymbol{q}$ and then fed into $C$ to filter out irrelevant documents.
  • CFS: To remove irrelevant dynamic-relevant documents, after retrieving $k_{1}$ documents, $<\boldsymbol{q},\boldsymbol{d}>$ pairs are matched one by one with the remaining documents for similarity. Simultaneously, they have been fed into $C$ for classification. If classified as negative, the process have been extended to the next document. Otherwise, positive instances will be included in the document set.
A.4 Baseline
We conduct a comprehensive comparison of our retrieval strategies against other RAG frameworks. In DR-RAG, we calculate the recall with different retrieval strategies and then evaluate the accuracy of the answers. Therefore, we select BM25 and SM methods Lewis et al. (2020); Chan et al. (2024) as baselines. Moreover, we choose self-RAG Asai et al. (2024) and Adaptive-RAG Jeong et al. (2024), which are effective RAG frameworks for multi-hop QA, to validate the performance of our DR-RAG. In addition, we add the experimental results of Non-retrieval, original RAG and multi-step approach Trivedi et al. (2023) to enrich our comparisons.
A.5 equation
A.5.1 QDC
$\begin{split}Context&=\{\}\\ \{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}&=\texttt{Retriever}(\boldsymbol{q})\\ Context&=Context\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}\\ \boldsymbol{q_{m}^{*}}&=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i})\\ \{\boldsymbol{d}^{\prime}_{1},\boldsymbol{d}^{\prime}_{2},\ldots,\boldsymbol{d}^{\prime}_{k}\}&=\texttt{Retriever}(\boldsymbol{q_{m}^{*}})\\ Context&=Context\cup\{\boldsymbol{d}^{\prime}_{j}\mid\boldsymbol{d}^{\prime}_{j}\not\in Context\text{ and }first\}\\ answer&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{q},Context))\end{split}$ (3)
A.5.2 CIS
$\text{Context}=\{\}$
$\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}=\texttt{Retriever}(\boldsymbol{q})$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{\frac{k}{2}}\}$
$\text{new}_{\text{query}}(i)=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i}),\quad i=1,2,\ldots,\frac{k}{2}$
$\{\boldsymbol{d}^{\prime}_{1},\boldsymbol{d}^{\prime}_{2},\ldots,\boldsymbol{d}^{\prime}_{k}\}=\texttt{Retriever}(\text{new}_{\text{query}}(i)),\quad i=1,2,\ldots,\frac{k}{2}$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}^{\prime}_{j}\mid\boldsymbol{d}^{\prime}_{j}\not\in\text{Context}\text{ and }\text{first}\},\quad i=1,\ldots,\frac{k}{2}$
$=\{\boldsymbol{d}_{i}\mid\exists j\neq i,C(\boldsymbol{d}_{i},\boldsymbol{d}_{j})\neq 0,\quad\boldsymbol{d}_{i},\boldsymbol{d}_{j}\in\text{Context}\}$ Context }
[…]
A.5.3 CFS
$\text{Context}=\{\}$
$\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{\frac{k}{2}}\}=\texttt{Retriever}(\boldsymbol{q})$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{\frac{k}{2}}\}$
$\text{new}_{\text{query}}(i)=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i}),\quad i=1,2,\ldots,\frac{k}{2}$
$\{\boldsymbol{d}^{\prime}_{1},\boldsymbol{d}^{\prime}_{2},\ldots,\boldsymbol{d}^{\prime}_{k}\}=\texttt{Retriever}(\text{new}_{\text{query}}(i)),\quad i=1,2,\ldots,\frac{k}{2}$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}^{\prime}_{j}\mid\boldsymbol{d}^{\prime}_{j}\not\in\text{Context}\text{ and }C(\boldsymbol{q},\boldsymbol{d}_{i},\boldsymbol{d}^{\prime}_{j})=1\text{ and }\text{first}\},\quad i=1,\ldots,\frac{k}{2}$
[…]
Table 7: Case of the classifier we train for select in CIS and CFS.
Dataset classified as positive classified as negative
musique
Query
In True Grit, who did the star play?
Document1
True Grit is a 1969 American western film. It is the first
film adaptation of Charles Portis’ 1968 novel of the same
name. The screenplay was written by Marguerite Roberts.
The film was directed by Henry Hathaway and starred
Kim Darby as Mattie Ross and John Wayne as U.S. Mar-
shal Rooster Cogburn. Wayne won his only Academy
Award for his performance in this film and reprised his
role for the 1975 sequel Rooster Cogburn.
Document2
In October 2015, TCM announced the launch of the TCM
Wineclub, in which they teamed up with Laithwaite to pro-
vide a line of mail-order wines from famous vineyards such
as famed writer-director-producer Francis Ford Coppola\’s
winery. Wines are available in 3 month subscriptions, and
can be selected as reds, whites, or a mixture of both. From
the wines chosen, TCM also includes recommended movies
to watch with each, such as a "True Grit" wine, to be paired
with the John Wayne film of the same name.
Query
In True Grit, who did the star play?
Document1
True Grit is a 1969 American western film. It is the first
film adaptation of Charles Portis’ 1968 novel of the same
name. The screenplay was written by Marguerite Roberts.
The film was directed by Henry Hathaway and starred
Kim Darby as Mattie Ross and John Wayne as U.S. Mar-
shal Rooster Cogburn. Wayne won his only Academy
Award for his performance in this film and reprised his
role for the 1975 sequel Rooster Cogburn.
Document2
The Iberian frog,Iberian stream frog or rana patilarga
("Rana iberica") is a species of frog in the family
Ranidae found in Portugal and Spain. Its natural ha-
bitats are rivers, mountain streams and swamps. It is
threatened by habitat loss, introduced species, climate
change, water contamination, and increased ultraviolet
radiation.
hotpotqa
Query
The 2000 British film Snatch was later adapted into a
television series for what streaming service?
Document1
Snatch is a British/American television series based
on the film of the same name, which debuted on Crackle
on 16 March 2017. The show has been renewed for a
second season.
Document2
Snatch (stylised as snatch.) is a 2000 British crime come-
dy film written and directed by Guy Ritchie, featuring an
ensemble cast. Set in the London criminal underworld,
the film contains two intertwined plots: one dealing with
the search for a stolen diamond, the other with a small-
time boxing promoter (Jason Statham) who finds himself
under the thumb of a ruthless gangster (Alan Ford) who
is ready and willing to have his subordinates carry out
severe and sadistic acts of violence.
Query
The 2000 British film Snatch was later adapted into a
television series for what streaming service?
Document1
Snatch is a British/American television series based
on the film of the same name, which debuted on Crackle
on 16 March 2017. The show has been renewed for a
second season.
Document2
Orange Is the New Black (sometimes abbreviated to
OITNB) is an American comedy-drama web television
series created by Jenji Kohan for Netflix. The series is
based on Piper Kerman\’s memoir, "" (2010), about her
experiences at FCI Danbury, a minimum-security federal
prison. "Orange Is the New Black" premiered on July 11,
2013 on the streaming service Netflix. In February 2016,
the series was renewed for a fifth, sixth, and seventh season.
The fifth season was released on June 9, 2017. The series
is produced by Tilted Productions in association with Lion-
sgate Television.
2wiki
Query
Where was the composer of film Love Story 1999 born?
Document1
Devanesan Chokkalingam, popularly known as Deva, is
an Indian film composer and singer. He has composed
songs and provided background music for Tamil, Telugu,
Malayalam and Kannada films in a career spanning about
20 years. Many know his gaana songs, written mostly
using Madras Tamil. He is known as the "Father of Gaana
Genre" in the Tamil film industry. Deva has composed
music for many films. He debuted as a film music director
in the film "Manasukkeththa Maharaasa" in 1989. In the
intervening years he has composed music for a total of
more than 400 films.
Document2
Love Story 1999 is a 1998 Indian Telugu-language romantic
comedy film directed by K. Raghavendra Rao. The film had
an ensemble cast starring Prabhu Deva, Vadde Naveen,
Ramya Krishna, Rambha and Laila in the lead roles.
Query
Where was the composer of film Love Story 1999 born?
Document1
Devanesan Chokkalingam, popularly known as Deva, is
an Indian film composer and singer. He has composed
songs and provided background music for Tamil, Telugu,
Malayalam and Kannada films in a career spanning about
20 years. Many know his gaana songs, written mostly
using Madras Tamil. He is known as the "Father of Gaana
Genre" in the Tamil film industry. Deva has composed
music for many films. He debuted as a film music director
in the film "Manasukkeththa Maharaasa" in 1989. In the
intervening years he has composed music for a total of
more than 400 films.
Document2
Tarcisio Fusco was an Italian composer of film scores. He
was the brother of the composer Giovanni Fusco and the
uncle of operatic soprano Cecilia Fusco.
Table 8: Cases the Query can be answered correctly in DR-RAG, and can not in Adaptive-RAG, where we present wrong answer in red and the right answer in blue.
Dataset Query and Answer Answer by Adaptive-RAG Answer by DR-RAG
MuSiQue
Query
What is the birth date of the By
Your Side performer?
True Answer
1 January 1986.
The performer of "By Your Side" is James
Cottriall. James Cottriall was born on 20
May 2011. So the answer is: 20 May 2011.
It is mentioned that James Cottriall was
born on 1 January 1986. Additionally, in
context, it is stated that "By Your Side"
is a song by James Cottriall.Putting this
information together, we can determine
that the birth date of James Cottriall, the
performer of the song "By Your Side," is
1 January 1986.
Answer: <1 January 1986>
hotpotqa
Query
Stephen susco wrote a movie
produced by which american
film actor?
True Answer
Do Kyung-soo
The member of the South Korean-Chinese
boy group EXO who stars in the upcoming
movie "The Underdog" is Sehun. So the
answer is: Sehun.
It is mentioned that Do Kyung-soo, also
known by his stage name D.O., is a mem-
ber of EXO and has starred in various
television dramas and movies. In context,
it is mentioned that the upcoming South
Korean animated film "The Underdog" stars
Do Kyung-soo. Therefore, based on the
information provided, Do Kyung-soo stars
in the upcoming movie "The Underdog".
Answer: <Do Kyung-soo>
2Wiki
Query
Where was the place of death
of the director of film The First
Day Of Freedom?
True Answer
Florida
The director of The First Day of Freedom is
Aleksander Ford. Aleksander Ford died in
Warsaw, Poland. So the answer is: Warsaw,
Poland.
In context, it is mentioned that Aleksander
Ford committed suicide in 1980 in Naples,
Florida, United States. Given the information
provided, the place of death of Aleksander
Ford, the director of the film "The First Day
of Freedom," was Naples, Florida, United
States.
Answer: <Naples, Florida, United States.>
Table 9: A case of our prompt provided to LLMs.
You are a reading comprehension expert, and you need to complete a reading comprehension task.
documents
Document 1:
Walk, Don’t Run is a 1966 Technicolor comedy film directed by Charles Walters and starring Cary Grant in his final film role,
Samantha Eggar, and Jim Hutton. The film is a remake of the 1943 film" The More the Merrier" and is set during the Olympic
Games
Document 2:
Douglas Sirk( born Hans Detlef Sierck; 26 April 1897 – 14 January 1987) was a German film director best known for his work
in Hollywood melodramas of the 1950s. Sirk started his career in Germany as a stage and screen director, but he left to Holly-
wood in 1937 because his Jewish wife was persecuted by the Nazis. In the 1950s, he achieved his greatest commercial success
with film melodramas like" Imitation of Life All That Heaven Allows Written on the WindMagnificent Obsession" and" A Time
to Love and a Time to Die". While those films were initially panned by critics as sentimental women’s pictures, they are today
widely regarded by film directors, critics and scholars as masterpieces. His work is seen as" critique of the bourgeoisie in general
and of 1950s America in particular", while painting a" compassionate portrait of characters trapped by social conditions". Beyond
the surface of the film, Sirk worked with complex mise en scenes and lush Technicolor colors to subtly underline his message.
Document 3:
The Mall, The Merrier is a 2019 Philippine musical family comedy film directed by Barry Gonzales, starring Vice Ganda and
Anne Curtis. The film is co-produced by Star Cinema and Viva Films under the working title" Momalland". The film premiered
in Philippine cinemas on December 25, 2019 as one of the official entries to the 2019 Metro Manila Film Festival. " The Mall,
The Merrier" marks the first on- screen collaboration between Anne Curtis and Vice Ganda, both of whom are regular hosts in
the noontime variety show" It’s Showtime".
Document 4:
Robert Wallace Russell( January 19, 1912 – February 11, 1992) was an American writer for movies, plays, and documentaries.
He was nominated for two Academy Awards for Best Writing, Original Story and Best Writing, Screenplay on the 1943 film"
The More the Merrier". He died in 1992 in New York City, shortly after his 80th birthday.
Document 5:
Sleep, My Love is a 1948 American film noir directed by Douglas Sirk and starring Claudette Colbert, Robert Cummings and
Don Ameche.
Document 6:
The More the Merrier is a 1943 American comedy film made by Columbia Pictures which makes fun of the housing shortage
during World War II, especially in Washington, D.C. The picture stars Jean Arthur, Joel McCrea and Charles Coburn. The movie
was directed by George Stevens. The film was written by Richard Flournoy, Lewis R. Foster, Frank Ross, and Robert Russell,
from" Two’s a Crowd", an original story by Garson Kanin( uncredited). This film was remade in 1966 as" Walk, Do n’t Run",
with Cary Grant, Samantha Eggar and Jim Hutton.
After reading the documents above, answering the following question. Reasoning step by step. At last, you should output the final
result via the following format:
Answer: <your answer based on the documents>;
Please answer the question directly.
question
Which film has the director who died later, The More The Merrier or Sleep, My Love?
Give your analysis process first, and then output your answer in a specified format.
Table 10: A case of documents retrieved by QDC and CFS, where the documents we need are in blue, and the top-$k$ is 4.
Query Documents retrieved by QDC Documents retrieved by CFS
Who is the spouse of the
Green performer?
Document 1:
Green is the fourth studio album by British pro-
gressive rock musician Steve Hillage. Written
in spring 1977 at the same time as his previous
album, the funk-inflected "Motivation Radio"
(1977), "Green" was originally going to be re-
leased as "The Green Album" as a companion
to "The Red Album" (the originally intended
name for "Motivation Radio"). However, this
plan was dropped and after a US tour in late
1977, "Green" was recorded alone, primarily
in Dorking, Surrey, and in London.
Document 2:
"Little Green" is a song composed and performed
by Joni Mitchell. It is the third track on her 1971
album "Blue".
Document 3:
The Main Attraction is an album by American
jazz guitarist Grant Green featuring performances
recorded in 1976 and released on the Kudu label.
Document 4:
Grant’s First Stand is the debut album by Ameri-
can jazz guitarist Grant Green featuring perfor-
mances by Green recorded and released on the
Blue Note label in 1961. Earlier recordings made
by Green for Blue Note were released as "First
Session" in 2001.
Document 1:
Green is the fourth studio album by British pro-
gressive rock musician Steve Hillage. Written
in spring 1977 at the same time as his previous
album, the funk-inflected "Motivation Radio"
(1977), "Green" was originally going to be re-
leased as "The Green Album" as a companion
to "The Red Album" (the originally intended
name for "Motivation Radio"). However, this
plan was dropped and after a US tour in late
1977, "Green" was recorded alone, primarily
in Dorking, Surrey, and in London.
Document 2:
"Little Green" is a song composed and performed
by Joni Mitchell. It is the third track on her 1971
album "Blue".
Document 3:
Miquette Giraudy (born 9 February 1953, Nice,
France) is a keyboard player and vocalist, best
known for her work in Gong and with her partner
Steve Hillage. She and Hillage currently form the
core of the ambient band System 7. In addition to
her performances in music, she has also worked
as an actress, film editor and writer. In each role,
she has used different stage names.
Table 11: A case of documents retrieved by QDC and CIS, where the documents we need are in blue, and the top-$k$ is 4.
Query Documents retrieved by QDC Documents retrieved by CIS
Who is the spouse of the
Green performer?
Document 1:
Caroline LeRoy Webster (September 28, 1797
in New York City – February 26, 1882) was
the second wife of 19th Century statesman
Daniel Webster. Her father was Herman LeRoy,
who was once head of the commercial house
of Leroy, Bayard, McKiven Co., a largetrading
company that operated in different partsof the
world. Her father was also the first Holland
Consul to the United States. Caroline’s mother
was Hannah Cornell, daughter of the last Royal
Attorney General of the State of North Carolina.
Caroline was a descendant of Thomas Cornell.
Document 2:
Pierre Paul Leroy-Beaulieu (9 December 1843
in Saumur – 9 December 1916 in Paris) was a
French economist, brother of Henri Jean Baptiste
Anatole Leroy-Beaulieu, born at Saumur, Maine
-et-Loire on 9 December 1843, and educated in
Paris at the Lycée Bonaparte and the École de
Droit. He afterwards studied at Bonn and Berlin,
and on his return to Paris began to write for "Le
Temps", "Revue nationale" and "Revue contem-
poraine".
Document 3:
Daniel Fletcher Webster, commonly known as
Fletcher Webster (July 25, 1813 in Portsmouth,
New Hampshire – August 30, 1862) was the son
of renowned politician Daniel Webster and Grace
Fletcher Webster. He was educated at Harvard
College. During his father’s first term as Secretary
of State, Fletcher served as Chief Clerk of the
United States State Department which, at the time,
was the second most powerful office in the State
Department. As Chief Clerk, he delivered the news
of President William Henry Harrison’s death to the
new President, John Tyler.
Document 4:
Leroy, also Leeroy, LeeRoy, Lee Roy, LeRoy
or Le Roy, is both a male given name in English -
speaking countries and a family name of French
origin. Leroy (lørwa) is one of the most common
surnames in northern France. As a surname it is
sometimes written Le Roy, as a translation of
Breton Ar Roue. It is an archaic spelling of le roi,
meaning “the king ”and is the equivalent of the
English surname King.
Document 1:
Daniel Fletcher Webster, commonly known as
Fletcher Webster (July 25, 1813 in Portsmouth,
New Hampshire – August 30, 1862) was the
son of renowned politician Daniel Webster and
Grace Fletcher Webster. He was educated at
Harvard College. During his father’s first term
as Secretary of State, Fletcher served as Chief
Clerk of the United States State Department
which, at the time, was the second most powerful
office in the State Department. As Chief Clerk,
he delivered the news of President William Henry
Harrison’s death to the new President, John Tyler.
Document 2:
Caroline LeRoy Webster (September 28, 1797
in New York City – February 26, 1882) was
the second wife of 19th Century statesman
Daniel Webster. Her father was Herman LeRoy,
who was once head of the commercial house
of Leroy, Bayard, McKiven Co., a largetrading
company that operated in different partsof the
world. Her father was also the first Holland
Consul to the United States. Caroline’s mother
was Hannah Cornell, daughter of the last Royal
Attorney General of the State of North Carolina.
Caroline was a descendant of Thomas Cornell.
Document 3:
Leroy, also Leeroy, LeeRoy,
Lee Roy, LeRoyor Le Roy, is both a male given
name in English - speaking countries and a family
name of French origin. Leroy (lørwa) is one of
the most common surnames in northern France.
As a surname it is sometimes written Le Roy, as
a translation of Breton Ar Roue. It is an archaic
spelling of le roi, meaning “the king ”and is the
equivalent of the English surname King.
Table 12: Results of 500 samples sampled on MuSiQue based on gpt-3.5-turbo and gpt-4-turbo.
top-$k$ LLMs EM F1 Acc
3 gpt-3.5-turbo 26.00 36.20 35.00
gpt-4-turbo 33.80 46.30 47.20
4 gpt-3.5-turbo 25.80 37.60 38.60
gpt-4-turbo 34.20 47.80 51.00
6 gpt-3.5-turbo 28.40 41.10 40.60
gpt-4-turbo 39.40 52.70 54.20